### v1

# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

Abstract
Unsupervised automatic speech recognition (ASR) aims to learn the mapping between the speech signal and its corresponding textual transcription without the supervision of paired speech-text data. A word/phoneme in the speech signal is represented by a segment of speech signal with variable length and unknown boundary, and this segmental structure makes learning the mapping between speech and text challenging, especially without paired data. In this paper, we propose Reborn, Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR. Reborn alternates between (1) training a segmentation model that predicts the boundaries of the segmental structures in speech signals and (2) training the phoneme prediction model, whose input is a segmental structure segmented by the segmentation model, to predict a phoneme transcription. Since supervised data for training the segmentation model is not available, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We conduct extensive experiments and find that under the same setting, Reborn outperforms all prior unsupervised ASR models on LibriSpeech, TIMIT, and five non-English languages in Multilingual LibriSpeech. We comprehensively analyze why the boundaries learned by Reborn improve the unsupervised ASR performance. [^2] Our code will be available soon.
Machine Learning, ICML
1 Introduction
Automatic speech recognition (ASR) systems convert speech signals into their transcription texts. Most state-of-the-art (SoTA) ASR systems are supervisedly trained with a large amount of labeled speech-text paired data (Chen et al., 2021; Radford et al., 2023). However, more than 80% of languages in the world have limited access to speech-text paired data (Pratap et al., 2023), and collecting such paired data requires intensive labor and costs, limiting the applicability of supervised ASR systems to these languages. This leads to significant efforts devoted to developing unsupervised ASR (UASR) systems, which aim to learn the mapping between speech signals and textual transcriptions (words or phonemes) without any speech-text pairs (Liu et al., 2018; Chen et al., 2019; Baevski et al., 2021; Liu et al., 2022; Gao et al., 2023).
UASR models learn to align the distribution of input speech signal and output text without paired data. Learning to match two distributions of sequences unsupervisedly has been extensively studied in unsupervised neural machine translation (NMT) (Artetxe et al., 2018), where the aim is to learn a neural network that can translate text in a source language to text in a target language without paired data (Lample et al., 2018a, b). Prior works show that adversarial training (Goodfellow et al., 2014) can be used to learn such mappings (Lample et al., 2018a, b), which employs a generator learning to translate and a discriminator learning to distinguish generated text from the real text in the target language.
Can we adopt such an adversarial training scheme for UASR? In unsupervised NMT, text can be easily tokenized into sub-word tokens, so unsupervised NMT only needs to learn the mapping between the source and target language’s token embeddings. However, in UASR, a phoneme or word is represented by a variable-length segment in the speech signal whose boundaries are unknown. Moreover, the length of a speech signal is much longer than the length of its textual transcription. The above characteristics of speech make learning the mapping between the segmental structures in speech and textual transcription challenging. Existing works in unsupervised ASR rely on heuristics or handcrafted rules to obtain the boundaries of the segmental structures (Liu et al., 2018; Yeh et al., 2018; Baevski et al., 2021). Yet, such handcrafted boundaries are often sub-optimal, bottlenecking the performance of UASR.
This paper focuses on obtaining better segmental boundaries to improve the mapping between segmental structure and textual transcription. We propose Reborn (Reinforcement-Learned Boundary Segmentation with Iterative Training), an unsupervised ASR framework with a segmentation model and a phoneme prediction model. The segmentation model determines segmental structure boundaries in speech signals, while the phoneme prediction model assigns a phoneme to each segmental structure. We use an iterative algorithm that alternates between two stages to train the segmentation model and the phoneme prediction model: The first stage trains the segmentation model. Since we do not have the ground truth segmental boundaries, we use reinforcement learning (RL) to provide unsupervised guidance for training the segmentation model. We experiment with various learning objectives and implement them as reward functions for learning the segmentation model. In the second stage, based on the boundaries predicted by the segmentation model learned in the previous stage, we use an adversarial loss similar to generative adversarial networks (GANs) (Goodfellow et al., 2014) to train the phoneme prediction model.
We conduct extensive experiments on LibriSpeech (Panayotov et al., 2015), TIMIT (Garofolo, 1993), and Multilingual LibriSpeech (MLS) (Pratap et al., 2020) to compare the phoneme error rate (PER) and word error rate (WER) with prior works. We show that our method outperforms all prior UASR methods on LibriSpeech, TIMIT, and five languages in MLS when using the same amount of training data. We perform thorough ablation studies to show that iterative training and the rewards we design are critical to the performance of Reborn. By analyzing the segmental structure obtained by our segmentation model, we find that the segmental structures are acoustic units smaller than phonemes, which helps the phoneme prediction model predict more accurate transcriptions.
2 Related Work: Unsupervised ASR (UASR)
An ASR model takes speech signals as the input and predicts the textual transcriptions. Unsupervised ASR aims to train an ASR model without access to paired speech-text data. Instead, the only available training data is unlabeled speech data and unlabeled text data, while the correspondence between speech and text is not available. In this paper, we follow prior UASR works (Liu et al., 2018; Baevski et al., 2021) to predict phoneme transcriptions from speech signals. Learning the mapping between speech signals and phoneme sequences can be formulated as a distribution matching problem, where we want to learn an ASR model whose output phoneme sequences match the distribution of real phoneme sequences.
Yeh et al. (2018) solve this distribution matching problem by an unsupervised loss function based on the empirical output distribution matching (ODM) (Liu et al., 2017), which guides the ASR model to output phoneme sequences that have statistic distributions close to real phoneme sequences.
Liu et al. (2018) propose to use the generative adversarial network (GAN) (Goodfellow et al., 2014) for UASR. GAN solves the distribution matching problem by training a generator whose output distribution resembles a target distribution, and a discriminator is trained to distinguish the outputs of the generator and the samples from the target distribution. When using GAN to train unsupervised ASR, the generator is the phoneme prediction model. The phoneme prediction model takes in speech features of the segmental structures in the speech signal, and outputs a phoneme transcription. Prior works rely on heuristics or separately trained unsupervised phoneme segmentation models to find the boundary of the segmental structures (Wang et al., 2018; Liu et al., 2018; Chen et al., 2019).
With the emergence of foundation models, wav2vec-U (Baevski et al., 2021) shows that UASR performance can greatly benefit from using speech foundation models as the feature extractor. However, wav2vec-U still relies on some heuristic feature preprocessing steps based on $k$-means to find the boundaries of the segmental structure in the speech signal. In our paper, we call the segmentation boundaries obtained by the feature preprocessing in wav2vec-U as ”$k$-means segmentation”. EURO (Gao et al., 2023) changes the decoding algorithm used to decode the phoneme prediction sequences and explores other speech foundation models for feature extraction, including HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022).
In these prior methods, the segmentation is either non-trainable or learned separately from the phoneme prediction model. Reborn differs from prior works by introducing a trainable segmentation model tailored for the phoneme prediction model, and the segmentation model and phoneme prediction model can be iteratively learned. wav2vec-U 2.0 (Liu et al., 2022) simplifies the feature preprocessing step in wav2vec-U and does not explicitly consider the segmentation structure in their model. Reborn has performance better than wav2vec-U 2.0 on almost all datasets, showing that learning the segmentation boundary for feature preprocessing is important to the performance of UASR.
Figure 1: (a) and (b): Reborn iterates between using RL to train the segmentation model and using adversarial training to train the phoneme prediction model. (c): An illustration of the segmentation post-processing. Given the original segmentation and the predicted phoneme sequence, we merge the segments that result in the same phoneme prediction into the same segment, yielding the post-processed boundary.
3 Method: Reborn
The difficulty of mapping speech signals to their corresponding phoneme transcriptions lies in the segmental structures in speech whose boundaries are unknown. To tackle this challenge, we propose Reborn, which trains a UASR system with unpaired speech and text data. Reborn contains a segmentation model and a phoneme prediction model. The segmentation model takes the speech feature as input and determines the boundaries of the segmental structures in the speech signal, and the phoneme prediction model predicts a phoneme for each segmental structure. We will use the terms segment and segmental structure interchangeably.
The overall training process of Reborn is outlined in Algorithm 1 and Figure 1. First, we initialize the phoneme prediction model using wav2vec-U (Section 3.3). Next, the training of Reborn iterates through two stages: Stage 1 (Figure 1(a), Section 3.1): Train the segmentation model to learn a better segmentation until the segmentation model converges while fixing the phoneme prediction model. Stage 2 (Figure 1(b), Section 3.2): Training the phoneme predictor based on the segment boundaries predicted by the segmentation model until the phoneme prediction model converges. The iterative process ceases when the UASR performance does not improve over the previous iteration. In UASR, we cannot use PER as a validation metric to select the best model or determine when to stop training. We use an unsupervised evaluation metric to select the best model, detailed in Appendix C.2.
3.1 Stage 1: Training the Segmentation Model
3.1.1   Segmentation Model
Given an input speech signal, we use a self-supervised feature extractor model to extract the speech features $X=[x_{1},\cdots,x_{T}]$ from the waveform. The segmentation model takes the speech features and predicts the boundary of the segmental structure in the speech features. For each feature $x_{i}$, the segmentation model assigns a binary value $\hat{b}_{i}$, indicating whether $x_{i}$ is the first frame of a segment. We split $X$ into segments based on $\hat{b}_{1:T}$ and mean-pool the features within the same segment. The mean-pooled features $[s_{1},\cdots,s_{T^{\prime}}]$ are forwarded through the phoneme prediction model to obtain the phoneme prediction $[y_{1},\cdots,y_{T^{\prime}}]$. The resulting phoneme prediction will be de-duplicated into $[y^{\prime}_{1},\cdots,y^{\prime}_{M}]$ by removing the same consecutive phoneme.
The segmentation model is a one-dimensional CNN, denoted by $\pi_{\theta}$ and parametrized by $\theta$. Given $[x_{1},\cdots,x_{T}]$, $\pi_{\theta}$ predicts a probability $\pi_{\theta}[i],$ $\forall i\in[1,\cdots,T]$. Let $\hat{B}_{i}\sim\text{Bernoulli}(\pi_{\theta}[i])$ be a Bernoulli random variable representing whether $x_{i}$ is the beginning of a segment. $\hat{B}_{i}=1$ means $x_{i}$ is the beginning of a segment; $\hat{B}_{i}=0$ means $x_{i}$ is not the beginning of a segment. During training, we sample $\hat{b}_{i}$ from $\text{Bernoulli}(\pi_{\theta}[i])$; during inference, we take $\hat{b}_{i}=1$ if $\pi_{\theta}[i]\geq 0.5$, otherwise $\hat{b}_{i}=0$.
3.1.2   Training the Segmentation Model with RL
To help the prediction of the phoneme prediction model, we want to train the segmentation model to capture the segmental structure in speech. However, the optimal segmentation boundary is not available for training. Recognizing that the segmentation quality directly affects the phoneme prediction of the phoneme prediction model, we estimate the quality of the phoneme prediction to guide the segmentation model training. The phoneme prediction model is from the previous Reborn iteration or initialized from wav2vec-U in the first Reborn iteration, and it is fixed when training the segmentation model. Since the effect of feature segmentation on phoneme prediction is non-differentiable, we employ RL to train the segmentation model. The estimated quality of the phoneme sequence serves as the RL reward.
We train the segmentation model using the policy gradient method (Sutton et al., 1999). For each utterance, we calculate an utterance-wise reward $R$ (defined in Eq. 4 in the next subsection) from the de-duplicated phoneme prediction to train the segmentation model. Based on the policy gradient method, the segmentation model $\pi_{\theta}$ is optimized using the following gradient: $\mathbb{E}\left[\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(\hat{b}_{t}|X)R\right]$, where $\mathbb{E}$ is taken over the training data distribution and approximated with the mean of a batch of training sequences.
3.1.3   Reward Designs
Given an utterance in the training set, the utterance-wise reward $R$ is the weighted sum of perplexity difference reward $R_{\text{ppl}}$, edit-distance reward $R_{\text{edit}}$, and length difference reward $R_{\text{len}}$.
We introduce some notations: Given an utterance, we use the currently trained segmentation model $\pi_{\theta}$ for segmentation and the phoneme prediction model trained in the last Reborn iteration to obtain a deduplicated phoneme prediction sequence, denoted by $Y^{\prime}_{\theta}$. For the same utterance, we can use the segmentation model trained in the previous Reborn iteration $\pi_{\theta-1}$ for features segmentation and the same phoneme prediction model to obtain another de-duplicated phoneme prediction sequence, denoted as $Y^{\prime}_{\theta-1}$. In the first Reborn iteration, $Y^{\prime}_{\theta-1}$ is the de-duplicated phoneme prediction from wav2vec-U.
Perplexity Difference Reward
The perplexity difference reward is designed to favor phoneme segmentation better than the segmentation learned in the previous iteration. Intuitively, a better phoneme segmentation prediction $\hat{b}_{1:T^{\prime}}$ should yield a more reasonable phoneme prediction $y_{1:T^{\prime}}$. We use perplexity (PPL), the negative likelihood of a phoneme sequence scored by a phoneme language model (LM), to evaluate how reasonable a phoneme sequence is. Perplexity measures how likely a phoneme sequence can be observed in the real world; a phoneme sequence with a lower perplexity means it is more likely to be observed in real-world phoneme datasets. We use KenLM (Heafield, 2011) as the phoneme LM, where the LM is trained with the unlabeled text corpora. To guide the segmentation model to generate a segmentation with a better phoneme prediction than the phoneme predictions obtained in the previous iteration, we define perplexity difference reward as follows:
$R_{\text{ppl}}=\text{PPL}_{\theta-1}-\text{PPL}_{\text{$\theta$}},$ (1)
where $\text{PPL}_{\theta-1}$ is the perplexity of $Y^{\prime}_{\theta-1}$ and $\text{PPL}_{\theta}$ is the perplexity of $Y^{\prime}_{\theta}$. The perplexity difference reward guides the segmentation model to produce segmentation that results in a phoneme sequence with lower perplexity.
When only using $R_{\text{ppl}}$, the segmentation model learns some segmentation method that leads to phoneme predictions with more $<$silence$>$ token, which has a lower perplexity but does not correspond to better phoneme prediction results. To avoid such unwanted behaviors, we design two regularization rewards, edit distance reward and length difference reward, to ensure that the policy learned by the segmentation model does not deviate too much from the segmentation learned in the previous iteration. The two regularization rewards punish the segmentation model if $Y^{\prime}_{\theta}$ is too different from $Y^{\prime}_{\theta-1}$. Crucially, regularization works only if the $Y^{\prime}_{\theta-1}$ in the first iteration is good enough. As previously mentioned, we ensure this by using the predictions from a trained wav2vec-U as $Y^{\prime}_{\theta-1}$ in the first iteration.
Edit Distance Reward
We use Levenshtein distance $d_{\text{Lev}}$ as the edit distance to calculate the difference between $Y^{\prime}_{\theta-1}$ and $Y^{\prime}_{\theta}$. We take the normalized negative edit distance as the reward, which is defined as follows:
$R_{\text{edit}}=-\frac{d_{\text{Lev}}(Y^{\prime}_{\theta-1},Y^{\prime}_{\theta})}{|Y^{\prime}_{\theta-1}|}$ (2)
Length Difference Reward
We use the length difference reward $R_{\text{len}}$ to guide the segmentation model to predict segmentation such that the length of $Y^{\prime}_{\theta}$ does not differ significantly from the length of $Y^{\prime}_{\theta-1}$. The length difference reward for an utterance is defined as:
$R_{\text{len}}=1-\frac{\left||Y^{\prime}_{\theta}|-|Y^{\prime}_{\theta-1}|\right|}{|Y^{\prime}_{\theta-1}|},$ (3)
where $|Y^{\prime}|$ is the length of $Y^{\prime}$.
Utterance-Wise Reward
The final utterance-wise reward $R$ is the weighted sum of $R_{\text{ppl}}$, $R_{\text{edit}}$, and $R_{\text{len}}$.
$R=c_{\text{ppl}}\cdot R_{\text{ppl}}+c_{\text{edit}}\cdot R_{\text{edit}}+c_{\text{len}}\cdot R_{\text{len}},$ (4)
where $c_{\text{ppl}}$, $c_{\text{edit}}$ and $c_{\text{len}}$ are the weighting coefficients. During training, $R_{\text{ppl}}$, $R_{\text{edit}}$ and $R_{\text{len}}$ are normalized within each training batch.Appendix C.2 discusses how the coefficients are determined without speech-text paired data.
3.1.4   Initializing $\pi_{\theta}$ with Behavior Cloning
Before training the segmentation model $\pi_{\theta}$ with RL, we initialize it with behavior cloning (BC). BC uses supervised training to train the segmentation model to predict the boundaries in the speech features. Given speech features $x_{1:T}$, the segmentation model is trained to predict the 0/1 boundary $\hat{b}_{i}$ by using some boundaries $b_{i}$ as the ground truth. The ground truth boundary is the post-processed boundary (Section 3.1.5) obtained using $\pi_{\theta-1}$, the segmentation model learned in the previous iteration. The BC process is a binary classification task that can be trained with cross-entropy loss. In the first Reborn iteration, we use the $k$-means segmentation boundary from wav2vec-U as the BC target.
3.1.5   Boundary Post-Processing
After training the segmentation model until convergence, we use the segmentation model to predict the boundaries of the whole dataset and post-process those boundaries. The post-processing step uses the phoneme prediction model trained in the previous iteration to refine the boundary predictions $\hat{b}_{i}$. Even if some consecutive speech features are split into two segments by the segmentation model, the mean-pooled features of the two segments may yield the same phoneme prediction. In this case, the two segments will be merged into one segment. An illustration of post-processing is shown in Figure 1(c). Post-processing differs from phoneme sequence de-duplication: post-processing modifies the segmentation prediction $\hat{b}_{i}$, while de-duplication modifies the phoneme sequence.
3.2 Stage 2: Training the Phoneme Prediction Model
The phoneme predictor takes the mean-pooled features of segmental structures $S=[s_{1},s_{2},...,s_{T^{\prime}}]$ and predicts a phoneme sequence $Y=[y_{1},y_{2},...,y_{T^{\prime}}]$. $S$ are obtained by mean-pooling the features $X$ in the same segmental structure based on $[\hat{b}^{{}^{\prime}}_{1},\cdots,\hat{b}^{{}^{\prime}}_{T^{\prime}}]$, where $\hat{b}^{{}^{\prime}}_{i}$ is the post-processed boundaries. We find that it is essential to perform boundary post-processing to stabilize the training in this stage.
After obtaining the post-processed phoneme boundaries, we train the phoneme predictor model using GAN training. The phoneme prediction model is the generator in GAN training. The generator aims to output phoneme predictions that look like real phoneme sequences to fool the discriminator. The discriminator takes in a phoneme sequence, which can be the output of the generator or a phoneme sequence from the unpaired text corpora, and the goal of the discriminator is to distinguish whether the input phoneme sequences are outputs of the generator. The generator and discriminator are updated using the loss in GAN training. In this stage, the parameters of the segmentation model are not updated.
3.3 Initialization of Reborn
In Stage 1, the segmentation model depends on the phoneme prediction when calculating the rewards. As a result, Reborn cannot work without properly initializing the phoneme prediction model. We use wav2vec-U to train a phoneme prediction model and use it as the initialization for the phoneme prediction model in Reborn. We briefly introduce wav2vec-U in Appendix A for interested readers.
Algorithm 1 Training process of Reborn
1:Initialization: Use wav2vec-U to obtain a phoneme predictor (§ 3.3)
2:while True do
3:     Stage 1: Initialize segmentation model with behavior cloning (§ 3.1.4) and then train using RL (§ 3.1)
4:     Post-process boundaries (§ 3.1.5)
5:     Stage 2: Train phoneme prediction model (§ 3.2)
6:     if no improvement (§ C.2then break
7:     end if
8:end while
4 Experiment Setup
4.1 Dataset
We use three datasets commonly used in ASR to evaluate the performance of Reborn.
LibriSpeech (Panayotov et al., 2015) is an English speech recognition corpora that contains 960 hours of training data. Following EURO (Gao et al., 2023), we use 100 hours of audio from the train-clean-100 set as the unlabeled speech data. The unlabeled text data does not overlap with the transcription of the unlabeled speech data.
TIMIT (Garofolo, 1993) is another English speech recognition with the human-labeled phoneme boundary. We follow the matched setting, which is more broadly used (Liu et al., 2018; Baevski et al., 2021; Gao et al., 2023), where the speech and text data come from the same set of utterances.
Multilingual LibriSpeech (MLS) (Pratap et al., 2020) is an ASR dataset including German (de), Dutch (nl), French (fr), Spanish (es), Italian (it), and Portuguese (pt). Following (Baevski et al., 2021), we use 100 hours of speech data for each language and use the LM data provided by the dataset as unpaired text.
4.2 Evaluation Metrics
We use phoneme error rate (PER) and word error rate (WER) to evaluate the ASR performance. If not specified, we use greedy decoding to obtain phoneme-level results. For decoding word-level outputs, we perform WFST decoding (Mohri, 1997; Mohri et al., 2002) using PyKaldi (Can et al., 2018). We leave the details in Appendix C.3.
4.3 Implementation Details
For English dataset, we use wav2vec 2.0 (Baevski et al., 2020) to extract speech features from speech signals; for MLS, we use XLSR-53 (Conneau et al., 2021) as the feature extractor. The segmentation model is a two-layer one-dimensional CNN model, detailed in Appendix C.1. For LibriSpeech and TIMIT, we train Reborn for two iterations. For MLS, we only train one iteration since we do not find the performance to improve in the second iteration.
5 Results
5.1 Main Results
We show the PER and WER of LibriSpeech, TIMIT, and MLS in Table 1, Table 2, and Table 3. We compare Reborn with several prior UASR works: wav2vec-U (Baevski et al., 2021), wav2vec-U 2.0 (Liu et al., 2022), and EURO (Gao et al., 2023). We have the following observations:
Table 1: PER/WER on LibriSpeech using 100 hours speech data. $\dagger$: Our reproduction. (wav2vec-U and wav2vec-U 2.0 only report results of using 960 hours of unlabeled speech).
Approach PER/WER (%) $\downarrow$ dev-clean dev-other test-clean test-other With Oracle Boundary Train from oracle 6.3/12.8 9.7/16.3 6.4/12.7 10.0/16.8 Baseline EURO (HuBERT) 15.2/23.1 20.7/29.3 15.1/22.8 21.1/29.8 ${}^{\dagger}$ 19.3/20.4 22.9/25.6 19.3/21.0 23.2/25.8 wav2vec-U ${}^{\dagger}$ 12.2/17.2 16.3/21.7 12.6/17.7 16.3/22.2 wav2vec-U 2.0 + HMM ST ${}^{\dagger}$ 10.0/15.4 13.1/19.0 10.3/16.0 13.1/19.6 Our Method Reborn 8.3/12.5 11.9/17.6 8.9/13.1 12.5/18.7 Reborn + HMM ST 5.2/9.3 8.5/13.5 5.4/9.6 8.5/13.7
Reborn significantly outperforms all prior methods on LibriSpeech
Our experiment on LibriSpeech follows Gao et al. (2023) to use the 100-hour training split of LibriSpeech as the unlabeled speech data. Compared with all the baselines in Table 1, Reborn achieves the lowest PER and WER. wav2vec-U and EURO both use heuristics and hand-crafted rules to obtain the segmental boundaries, while wav2vec-U 2.0 removes the feature segmentation steps. The superior performance of Reborn illustrates the importance of learning the segmental boundaries in UASR. Notably, without using HMM self-training, Reborn already has PER/WER lower than wav2vec-U 2.0 with HMM self-training, which is the prior SoTA method on LibriSpeech.
Self-training further improves PER/WER of Reborn
Self-training uses the phoneme predictor’s prediction as the pseudo-label and trains a new phoneme prediction model using the pseudo-label as the training data. It is commonly used in UASR to boost the performance (Baevski et al., 2021; Liu et al., 2022). In Table 1, we show that Reborn can be integrated with HMM self-training to further lower the PER/WER. We reach a new SoTA on LibriSpeech under the setting of 100-hour speech data, outperforming the prior SoTA in PER and WER by 5% and 6%, respectively. Surprisingly, Reborn with HMM self-training outperforms training wav2vec-U with the oracle boundary. This shows that Reborn can be combined with self-training to improve performance effectively. Due to limited computation resources, we only conduct self-training on LibriSpeech.
Table 2: PER results on TIMIT. The cross-mark () in the greedy-decoding column indicates that an additional LM is used during decoding.
Approach Greedy decoding PER (%) core-dev core-test all-test (a) Train from oracle 9.1 10.3 9.2 (b) wav2vec-U (reproduced) 20.2 22.2 20.3 (c) wav2vec-U+WFST 17.1 17.8 16.8 (d) EURO (wav2vec 2.0) 18.5 19.8 - (e) EURO (WavLM) 14.3 14.6 - (f) Reborn 12.4 13.5 12.4
Reborn outperforms all prior UASR methods on TIMIT
Table 2 shows the PER of TIMIT, and we again find that Reborn achieves the lowest PER compared with all prior UASR methods. This indicates that Reborn not only works on large datasets like LibriSpeech but also works on small datasets like TIMIT, which only contains about 3 hours of audio data for training. Even using greedy decoding only, Reborn outperforms the prior best-performing UASR model (row (e)), which relies on prefix decoding. EURO shows that replacing the feature extractor with WavLM (Chen et al., 2022) (row (e)) outperforms using wav2vec 2.0 (row (d)) as the speech feature extractor. Since Reborn using wav2vec 2.0 already outperforms EURO with WavLM, we leave changing the feature extractor in Reborn as future work.
Table 3: PER/WER on MLS. $\dagger$: Results from Baevski et al. (2021). ***: Our reproduction of wav2vec-U, used as the initialization of the phoneme prediction model in Reborn. $\ddagger$: Results from Liu et al. (2022).
Approach WER (%) $\downarrow$ de nl fr es it pt Avg. ${}^{\dagger}$ 32.5 40.2 39.8 33.3 58.1 59.8 44.0 $*{}^{*}$ 33.9 38.1 37.7 33.1 51.8 59.4 42.3 wav2vec-U ${}^{\ddagger}$ 23.5 35.1 35.7 25.8 46.9 48.5 35.9 Reborn 20.9 26.9 28.2 24.7 39.9 51.5 32.0
Figure 2: Comparing Reborn segmentation (a) with oracle boundaries (c) on an example (”It’s healthier to cook…”) from TIMIT test set. The ground truth phoneme transcription for each segment is given in (b). Silences are denoted with a short dash (-). The red crosses indicate wrong phonemes w.r.t. the ground truth; while the green ticks indicate the corrected ones. The dotted lines illustrate the delayed phenomenon as mentioned in Section 5.2.
Reborn performs well on languages other than English
The results on MLS are presented in Table 3. We find out that a single iteration is enough for performance convergence on MLS. Recall that Reborn is initialized from wav2vec-U, but Reborn outperforms wav2vec-U by a large margin in all six languages. This large performance gap underlines the importance of learning the boundaries of the segmental boundaries. Moreover, Reborn outperforms wav2vec-U 2.0 except for Portuguese, and the average WER on the six languages is 3.9% lower than wav2vec-U 2.0.
5.2 Boundary Analysis
The core of Reborn is the segmentation model that learns the segmental structure’s boundaries. In this section, we look at how the quality of segmental boundaries affects the performance of phoneme prediction. We compare the boundaries obtained by five methods: (a) the $k$-means segmentation boundary in wav2vec-U, (b) the oracle phoneme boundaries obtained by forced alignment (McAuliffe et al., 2017), which requires paired speech-text data to learn the alignment; (c) the phoneme boundary obtained by the SoTA unsupervised phoneme segmentation method (Strgar & Harwath, 2022); (d) the boundaries learned by the segmentation model of Reborn in the first iteration before post-processing and (e) after post-processing.
Table 4: Boundary evaluation results of different segmentation methods on LibriSpeech test-clean split. The second-last column (Freq.) is the number of segments per second. All the methods shares the same phoneme prediction model trained with the $k$-means segmentation of wav2vec-U.
Boundary method Prec. Rec. F1 Freq. PER (%) (a) Oracle 1.0 1.0 1.0 11.89 13.8 (b) $k$-means 0.64 0.77 0.70 14.27 19.3 (c) Strgar & Harwath (2022) 0.75 0.75 0.75 12.26 23.2 (d) Reborn (no post-processing) 0.57 0.78 0.65 16.29 12.9 (e)  + post-processing 0.73 0.71 0.72 11.63 - (f)  + left-shift 1 frame 0.76 0.74 0.75 11.63 -
First, we focus on the PER in Table 4 when using different segmentation methods. A segmentation method segments the speech feature into segments, and the segment-wise mean-pooled features will be used as input to the phoneme prediction model. In Table 4, all the segmentation methods share the same phoneme prediction model: the phoneme prediction model trained using the $k$-means segmentation of wav2vec-U. We find that simply replacing the $k$-means boundary with the oracle phoneme boundary reduces the PER by 5.5%, showing the imperfect heuristic segmentation is the bottleneck of UASR. Next, even when using the boundary predicted by the SoTA unsupervised phoneme segmentation (Strgar & Harwath, 2022), the PER does not reduce. Conversely, the boundary learned by Reborn achieves the lowest PER. This is because the boundaries learned by Reborn are tailored for the phoneme prediction model.
Next, we discuss the phoneme boundary quality in Table 4. The phoneme boundary quality is evaluated with boundary precision, recall, and F1 with a 20ms tolerance window, following the harsh scheme from Strgar & Harwath (2022). We also report the average number of segments per second (Freq. in Table 4). Interestingly, the boundary F1s of the oracle boundary (row (a)) and Strgar & Harwath (2022) (row (c)) are much higher than the boundary F1 of Reborn (row (d)), but the PER of Reborn is much better. This is because Reborn’s segmentation model predicts more segments than the number of phonemes in the utterance (before post-processing), indicating that it learns some segmental structures smaller than the phonemes. This can be seen from the lower precision (0.57) and higher frequency (16.29) compared to oracle boundaries. However, overly segmenting the speech feature is not problematic because even if the consecutive speech features of the same phonemes are split into two segments by the segmentation model, as long as they have the same phoneme prediction, the duplicated phoneme prediction will be removed in the de-duplication step, and thus not affecting the PER.
While Reborn boundaries initially have low precision, with post-processing, boundary precision can be improved by 0.16. Additionally, after applying boundary post-processing, the number of segmental structures is closer to the oracle. In contrast, we find that applying post-processing does not improve the boundary F1 for all other unsupervised segmentation methods listed in Table 4.
We show an example utterance in Figure 2 with the oracle boundary and Reborn boundary (before post-processing). One can observe that Reborn boundaries produce more segments than the oracle boundaries. We also observe an interesting boundary delay phenomenon: the phoneme boundaries predicted by Reborn sometimes lag behind the oracle phoneme boundaries (dotted lines in Figure 2). This observation motivates us to left-shift the post-processed Reborn boundary by one frame, and we find that the resulting boundary F1 is further improved and matches the boundary F1 of the unsupervised phoneme segmentation SoTA method (Strgar & Harwath, 2022) (row (f) v.s. row (c)).
5.3 Ablation Study
5.3.1   Reward Design and Behavior Cloning
Table 5: We compare different reward functions and the effect of BC initialization with first iteration results on LibriSpeech. Row (c) is Reborn.
Ablation Stage 1 PER (%) Stage 2 PER (%) Silence Ratio dev-clean test-clean dev-clean test-clean (a). $R_{\text{ppl}}$ 14.0 14.7 12.2 12.9 0.13 (b). $R_{\text{ppl}}+R_{\text{edit}}$ 13.3 13.8 11.7 12.1 0.07 (c). $R_{\text{ppl}}+R_{\text{edit}}+R_{\text{len}}$ 12.7 12.9 11.7 11.9 0.05 (d). Reborn w/o BC 14.8 14.9 14.1 14.2 0.05
We verify the effectiveness of our three proposed reward functions and behavior cloning initialization in Table 5. While the perplexity difference reward alone is sufficient to guide the segmentation model, adding edit distance and length difference rewards improves performance further. As previously mentioned, only using the perplexity difference reward will make the segmentation model predict more segments that lead to phoneme prediction sequences with more $<$silence$>$ tokens, which have a lower PPL but do not correspond to better phoneme prediction results. To demonstrate this, we show the silence ratio in the de-duplicated phoneme prediction sequence per utterance. We find that the silence ratio when only using $R_{\text{ppl}}$ (row (a)) is 0.13, which is much higher than the silence ratio of ground truth phoneme transcription sequences, which is 0.07. Adding $R_{\text{edit}}$ (row (b)) and $R_{\text{len}}$ (row (c)) makes the silent ratio closer to the ground truth silence ratio.
We also find that removing BC (row (d)) worsens PER compared with using BC (row (c)). Additionally, BC greatly speeds up RL training convergence, as shown in Figure 4 (detailed in Appendix B).
Figure 3: PER of each stage during Reborn’s two-stage iterative training on the test-clean split of LibriSpeech. St.: stage; w2vu: wav2vec-U
5.3.2   Iterative Training
In Figure 3, we show the PER on LibriSpeech test-clean split during the iterative training process of Reborn. We observe that after training the segmentation model in the first iteration, the PER drastically drops by 6.4% compared with wav2vec-U, which is used as the initialization of the phoneme prediction model in iteration one. This shows that the quality of the segmental structure’s boundaries is important to the performance of UASR. Afterward, the PER of Reborn decreases steadily across iterative training, showing that training the segmentation model and the phoneme prediction model is critical to improving the performance of Reborn. We find that the PER does not drop significantly after the fifth iteration.
6 Conclusion
In unsupervised ASR, learning the mapping between the segmental structures in speech and their corresponding transcription is challenging, especially without paired data. To tackle this, we propose Reborn, an iterative training method for unsupervised ASR. Reborn iteratively trains the segmentation model and the phoneme prediction model to improve the ASR performance. The segmentation model is trained using RL with carefully designed rewards to guide the segmentation model to find the boundaries of the segmental structures in speech signals. Extensive experiments on three datasets spanning seven languages show that Reborn outperforms prior best-performing unsupervised ASR models in all datasets except one language in MLS. We conduct comprehensive ablation studies to show that each component in Reborn is critical to the performance. We also explain the effectiveness of Reborn by analyzing its segmentation patterns and find that Reborn tends to produce segmental structures smaller than phonemes, which helps the generators to better predict the phoneme transcription.
Impact Statements
Unsupervised automatic speech recognition (UASR) - predicting the textual transcriptions for the given speech signals using only unpaired speech and text data - is highly attractive. The existence of thousands of low-resourced languages spoken globally is the major reason. This paper presents Reborn, a novel unsupervised ASR training algorithm. We foresee that Reborn has a great potential to make low-resource languages more accessible to train ASR systems, making speech technology more accessible and boundaryless.
However, our training method may reinforce the bias in the dataset. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A wav2vec-U
wav2vec-U (Baevski et al., 2021) is a popular UASR framework whose variants (Gao et al., 2023; Liu et al., 2022) have achieved SoTA performance on multiple UASR benchmarks. Following prior works (Liu et al., 2018), wav2vec-U also uses adversarial training to train the phoneme prediction model. The wav2vec-U framework takes the raw waveform as the input and predicts phoneme transcription via three stages: (1) the feature extraction stage, (2) the feature preprocessing stage, and (3) the phoneme prediction stage.
Stage (1) extracts speech features using a self-supervised speech foundation model, wav2vec 2.0 (Baevski et al., 2020). For a spoken utterance, we denote the extracted feature as $X=[x_{1},x_{2},...,x_{T}]\in\mathbb{R}^{T\times d}$, where $T$ is the time dimension and $d$ is the feature dimension. Stage (2), the feature preprocessing stage, uses $k$-means clustering and additional heuristics, namely, adjacent pooling, to define the boundary of the segmental structures in the speech features. The features in the same segmental structure are mean-pooled into one feature embedding to reduce sequence length to match phoneme sequences better. The pooled features’ dimension is reduced using PCA. We denote the output from the feature preprocessing stage as $S=[s_{1},s_{2},...,s_{T^{\prime}}]\in\mathbb{R}^{T^{\prime}\times d^{\prime}}$. In Stage (3), the phoneme prediction model takes the pooled speech feature as input and predicts the phoneme sequence $Y=[y_{1},y_{2},...,y_{T^{\prime}}]$. Last, a de-duplication step removes the consecutive duplicate phoneme predictions from $Y$, resulting in a shortened phoneme sequence $Y^{\prime}=[y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{M}]$ without consecutive duplicate phonemes. $Y^{\prime}$ will be used as the input to fool the discriminator in adversarial training, which loss is defined as follows:
$\mathcal{L}_{GAN}=$ $\min_{G}\max_{D}\underset{Z\sim\mathcal{Z}}{\mathbb{E}}[\log(\mathcal{D}(Z))]+\underset{X\sim\mathcal{X}}{\mathbb{E}}[\log(1-\mathcal{D}(\mathcal{G}(X)))]$ (5)
where $Z$ is a phoneme sequence sampled from the text corpora $\mathcal{Z}$ and $X$ is the speech signal from the speech corpora $\mathcal{X}$. Besides the min-max objective in the adversarial training, wav2vec-U (Baevski et al., 2021) also adopts the three regularized objectives for better convergence and performance. First, the gradient penalty $\mathcal{L}_{gp}$ (Gulrajani et al., 2017) is applied on the discriminator $\mathcal{D}$ to stabilize training, where the input of $\mathcal{D}$ is the linear combination of a generated sample $Y$ and a real sample $Z$.
$\mathcal{L}_{gp}=\underset{Y,Z,\alpha\sim U(0,1)}{\mathbb{E}}\left[\left(\left\|\nabla\mathcal{D}(\alpha Y+(1-\alpha)Z)\right\|_{2}-1\right)^{2}\right]$ (6)
Next, the smoothness penalty $\mathcal{L}_{sp}$ is computed across consecutive segments of the generated sequence, promoting the generation of closely aligned neighboring outputs by the model.
$\mathcal{L}_{sp}=\sum_{(y_{t},y_{t+1})\in Y}\left\|y_{t+1}-y_{t}\right\|^{2}$ (7)
Furthermore, the phoneme diversity loss $\mathcal{L}_{pd}$ is applied to encourage higher vocabulary usage of the generated sequences. As shown in Equation 8, the entropy of the averaged softmax distribution of the generator outputs ($H_{\mathcal{G}}(\mathcal{G}(S))$) is maximized over the phoneme vocabulary across a batch B.
$\mathcal{L}_{pd}=\frac{1}{|B|}\sum_{S\in B}-H_{\mathcal{G}}(\mathcal{G}(S))$ (8)
Finally, we describe the overall training objectives $\mathcal{L}$ by weighting with coefficients $\lambda$, $\gamma$, and $\eta$, respectively.
$\mathcal{L}=\mathcal{L}_{GAN}-\lambda\mathcal{L}_{gp}+\gamma\mathcal{L}_{sp}+\eta\mathcal{L}_{pd}$ (9)
In this work, we utilize the same training objective from wav2vec-U to perform GAN training (§ 3.2)
Appendix B Impact of Behavior Cloning on Model Convergence
Figure 4: Training Dynamics of BC Pre-Trained vs. Scratch Models. This figure depicts the evolution of Phoneme Error Rate (PER) over training epochs, tested on the test-clean set, to showcase the enhanced convergence speed and efficiency achieved with BC pre-training.
In this ablation study, we rigorously examine the influence of Behavior Cloning (BC) pre-training on the convergence dynamics and overall efficacy of our segmenter model. Our empirical findings reveal a marked improvement in sample efficiency for models subjected to BC pre-training, which is a critical aspect in enhancing model performance. Notably, these models demonstrate a notably faster convergence rate as compared to their counterparts trained from scratch.
As depicted in Figure 4, the training trajectories of the models, with and without BC pre-training, are contrasted. This visualization distinctly underscores the difference in PER across various training epochs. The curves elucidate not only the expedited convergence but also the superior performance attainable with BC pre-trained models.
It is observed that segmenters solely pre-trained through BC do not exhibit a substantial improvement in PER initially. However, this pre-training imbues the model with a more refined capability for boundary prediction. In subsequent RL training phases, this enhanced boundary prediction proficiency translates into a swifter convergence trajectory, culminating in superior final performance metrics. This phenomenon highlights the nuanced yet significant role of BC pre-training in the context of unsupervised automatic speech recognition, underscoring its potential as a pivotal step in the model optimization process.
Appendix C Implementation Details
C.1 Segmentaiton model
We parametrize the segmentation model $\pi_{\theta}$ using a two-layered 1-dimensional convolutional neural network (CNN). This choice of architecture allows us to efficiently extract relevant information from adjacent speech features, making it particularly well-suited for our segmentation task. The three convolutional layers in our CNN have different kernel sizes, namely 7 and 3. This is a lightweight model, making it computationally efficient for applications.
C.2 Hyperparameter Searches
We search the optimal hyperparameters of our phoneme predictor and segmentation models with unsupervised validation metrics. For the phoneme predictor, we directly adopt the search space of the hyperparameters and the unsupervised validation metric from wav2vec-U (Baevski et al., 2021).We search the hyperparameters indicated as $\lambda$, $\gamma$, and $\eta$ in Appendix A only during the initialization stage of the phoneme predictor (§ 3.3). We directly adopt the same hyperparameters n the following iterations of the phoneme predictor GAN training. As for the segmentation model, we take the idea from Baevski et al. (2021) and derive a similar unsupervised metric.
C.2.1   Unsupervised Validation Metric for Segmentation Model Selection
Each segmentation model, denoted as $\pi_{\theta}$, undergoes unsupervised performance evaluation by predicting the transcriptions of the validation set. The phoneme predictions generated by the model are processed to form a de-duplicated set, denoted as $Y^{\prime}_{\theta}=[y^{\prime}_{1,\theta},\cdots,y^{\prime}_{M,\theta}]$, where consecutive duplicate phonemes are removed.
The unsupervised metric places significant emphasis on two key aspects: Language Model (LM) negative log-likelihood (NLL) and vocabulary usage. Vocabulary usage is calculated as $U(Y^{\prime}_{\theta})=\frac{1}{|V|}\sum_{v\in V}\mathbb{I}(v\in Y^{\prime}_{\theta})$, where $V$ denotes the entire vocabulary set. The validation metric seeks to minimize the NLL while simultaneously maximizing vocabulary usage. The NLL metric reflects the model’s proficiency in generating sensible sentences consistent with the dataset distribution, while vocabulary usage quantifies the diversity of phonemes employed by the model. This combined approach ensures that our unsupervised ASR system’s predictions align with the expected LM distribution and exhibit a wide linguistic range, thus mitigating the risk of favoring repetitive, high-probability sentences.
Additionally, the lengths of transcriptions are taken into account. We employ NLL without length normalization, thereby favoring predictions that score high under the language model but without an undue preference for overly lengthy transcriptions.
The optimization objective is formally expressed as:
$\theta^{*}=\arg\min_{\theta}\left(\frac{{-\sum_{i=1}^{M}\text{ log }p_{LM}(y^{\prime}_{i,\theta})}}{{U(Y^{\prime}_{\theta})}}\right)$ (10)
captures these considerations and serves as a comprehensive measure for the performance and generalization capabilities of our unsupervised ASR model.
C.2.2   Training Details
For the training of our segmentation model, we initially employ behavioral cloning, followed by updates using policy gradient methods. In the BC phase, we configure our model with a batch size of 128 and a learning rate of 0.0005, training for 20 epochs. We also assign cross-entropy weights of [1, 5] for boundary classification, to address the imbalance where the class labeled 00 significantly outnumbers the class labeled $1$.
During the policy gradient update phase, the model is trained with the same batch size of 128, but with a reduced learning rate of 0.0001 and a weight decay of 0.0001. We utilize the AdamW optimizer in conjunction with the CosineAnnealingLR scheduler. For the LS and MLS datasets, the training proceeds for 40 epochs, while for the TIMIT dataset, owing to its smaller size, we extend the training to 500 epochs to ensure convergence.
To refine the performance of our reinforcement learning framework, we have tuned the reward coefficients, leveraging the validation metric (eq. 10), to secure optimal outcomes across varied datasets. Specifically, the coefficient for the PPL difference reward, $c_{\text{ppl}}$, is fixed at 1.0 to prioritize the minimization of perplexity in generated sequences. The coefficient for the edit-distance reward, $c_{\text{edit}}$, is set at 0.2. And the coefficient for the length difference reward, $c_{\text{len}}$, is selected from a predefined set of values within the range [0.0, 0.2, 0.4, 0.6, 0.8]. These values are carefully chosen to balance the importance of length consistency with other factors, with the optimal reward configuration for each dataset detailed in Table 6.
It is noteworthy that, as the iterations progress, the length constraints become less critical. This observation suggests that the model gradually learns to generate sequences of appropriate length, reducing the need for explicit length-based adjustments.
Table 6: Best reward configurations obtained through hyperparameter searches on each dataset.
Dataset LibriSpeech TIMIT MLS iter. 1 iter. 2-5 iter. 1-3 de nl fr es it pt $c_{\text{ppl}}$ 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $c_{\text{edit}}$ 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 $c_{\text{len}}$ 0.2 0.0 0.0 0.4 0.2 0.8 0.4 0.6 0.4
C.3 WFST Decoding for Obtaining Word-Level Outputs
In Reborn, phoneme-level outputs can be generated directly with greedy decoding. However, obtaining word-level outputs requires additional modules such as WFST or external language models (Baevski et al., 2021; Liu et al., 2017; Gao et al., 2023). In this work, we incorporate Reborn with WFST to generate word-level transcriptions. Our decoding procedure mainly follows Baevski et al. (2021), which builds the WFST using PyKaldi (Can et al., 2018), and additional self-loops are introduced to mimic the CTC behavior. Moreover, two hyperparameters: the acoustic scale $a$, and the additive blank weight $v$ are added during the WFST decoding process. In wav2vec-U, the search interval of $a$ is $[0,8]$ and $v$ is tuned within $[-3,8]$. As for Reborn, we find that using a much smaller search space for both of the hyperparameters is enough for obtaining reasonable outputs. More specifically, We tune the acoustic scale in $[0.4,1.2]$ with step $0.1$ and the blank weight in $[-5,2]$ with step $1$.
It is worth mentioned that we utilize the original boundary generated by the segmentation model instead of the post-processed one (§ 3.1.5) for WFST decoding. We find that the increased amount of segments directly leads to longer generator outputs, which is beneficial for WFST to generate higher quality outputs. The phenomenon is also discovered by the prior works (Baevski et al., 2021; Liu et al., 2022). In wav2vec-U (Baevski et al., 2021), the segmental feature used for WFST decoding is the one before adjacent pooling; while in wav2vec-U 2.0 (Liu et al., 2022), they simply lowers the decoding stride to generate more lengthy outputs for WFST decoding.